Example 3 A particle travels along the path y-2x2. At a given instant, x-1 m and dx/dt - 2 m/s which is a constant value. Find the magnitude of velocity at this instant. (Answer: v-8.25 m/s)
Answer #1

Y = 2x^2

On differentiate , V = dy/dt = 4x.dx/dt

=) V = 4×1×2 = 8 m/s

Velocity V = 8m/s

I think answer should be 8m/s ... I think there may be any mistake in given answer. Please comment below in case if any query.